OpenStudy (antonio_xx2):

Match each set of measures of a and b with the measure of c that gives a volume of forty-eight cubic inches. a=4inches,b=3inches a=4inches,b=4inches a=8inches,b=inches a=6inches,b=4inches options are A)c=6inches B)c=2inches C)c=8iches D)c=4inches

hartnn (hartnn):

V= abc =48 so c= 48/(ab) plug in a and b values here, you will get corresponding c values :)

hartnn (hartnn):

can you post what you tried ?

OpenStudy (antonio_xx2):

sure

OpenStudy (antonio_xx2):

4*4=16. 48/16=3 and there is no 3

hartnn (hartnn):

a=4inches,b=3inches , c =4 a=6inches,b=4inches, c =2 a=4inches,b=4inches, c =3, not in choices a=8inches,b=inches ...idk what b is ...

OpenStudy (antonio_xx2):

b is wut

hartnn (hartnn):

thats what i am asking you posted a=8inches,b=inches how would i know the 'b' value ?

OpenStudy (antonio_xx2):

ow sorry its 6

hartnn (hartnn):

a=8inches,b=6 inches, c =1 not in choices...

OpenStudy (antonio_xx2):

yes?

hartnn (hartnn):

so either question is incorrect, or the choices

OpenStudy (antonio_xx2):

i think so

hartnn (hartnn):

if the volume was 96, then all choices would have been correct, a=4inches,b=3inches , c =8 a=6inches,b=4inches, c =4 a=4inches,b=4inches, c =6, a=8inches,b=6inches, c=2.

OpenStudy (antonio_xx2):

@AccidentalAiChan

OpenStudy (antonio_xx2):

@esshotwired

OpenStudy (antonio_xx2):

@Shallow00

OpenStudy (antonio_xx2):

@Shallow00